Fermi National Laboratory


Accelerator Update for May 6, 2005 - May 9, 2005

Friday May 6
The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, and with Main Injector (MI) BPM work in progress.

Operations put the Tevatron (TeV) into shot setup at 8:42 AM.

At 9:28 AM, Operations reported that beam was not making it into the Booster. Operators investigated and discovered a problem with the 400 MeV area chopper. An expert had the chopper back online by 10:24 AM.

At 11:46 AM, Operations reported that the first transfer of antiprotons was lost in MI. Garbage in the MECAR program caused the loss.

At 12:23 PM, Operations reported more problems with MECAR. An expert investigated.

Operations established store 4131 at 1:20 PM with an initial luminosity of 53.54E30.

Pbar resumed taking beam at 2:01 PM.

Operations established beam to the Meson MCenter experiment and to NuMI at 2:06 PM.

At 3:16 PM, Operations reported beam problems between the Pbar AP1 line and the P2 transfer line. Operators and experts investigated and at 4:30 PM found a problem with BV706.

Operations established beam to MiniBooNE at 5:12 PM, but beam problems at 5:23 PM were keeping them off.

At 6:48 PM, a Linac RF station (KRF2) tripped off and wouldn't reset. Operators called in an expert who had the station back on by 7:01 PM.

At 7:50 PM, Operations reported that beam stopped making it to the Pbar target. Operators called in an expert.

Operations put the Recycler into shot setup to stash antiprotons at 8:30 PM.

At 9:46 PM, Operations reported that the MIRF system went down on the third transfer of antiprotons. The antiprotons made it to the Recycler.

At 9:57 PM, Operations reported that the MiniBooNE beam was tripping off every few seconds due to beam problems. NuMI was also having beam problems. Operators investigated and discovered the problem with this and beam to the Pbar target was due to the Booster chopper. An Operator called in an expert.

Stashing the Recycler ended at 10:13 PM.

Operations tried to put the Pbar into stacking mode at 10:52 PM, but septa problems (I:SEP2V) delayed this. Operators called in experts who found a bad SCR at 11:14 PM.

Saturday May 7
The midnight shift began with Operations monitoring store 4131, the stash, and the Pbar stack, with Accumulator septa work continuing, and with MiniBooNE, NuMI, and MCenter taking beam.

Pbar resumed stacking at 2:01 AM.

At 2:09 AM, Operations reported more Booster chopper problems.

At 2:23 AM, septa experts reported they found and replaced a bad SCR firing module.

At 4:13 AM, the D-Zero (D0) experiment reported seeing a proton beam problem around 2 AM.

At 7 AM, a Booster RF station (BRF16) tripped off. It reset okay.

At 8:23 AM, BRF16 tripped off again. It reset okay.

At 8:48 AM, BRF15 tripped off. It reset okay.

At 10:28 AM, BRF16 tripped off for the first of two times in fourteen minutes. It reset okay both times.

Operations terminated store 4131 at 12:04 PM and put the TeV into shot setup at 12:22 PM.

Operation established store 4132 at 2:52 PM with an initial luminosity of 78.19E30. Pbar resumed stacking.

At 3:51 PM, NuMI lost its beam permit for the first of two times in forty minutes. It reset okay for the first trip, but the second trip took about thirty minutes to investigate and reset the beam.

At 5:32 PM, Operations reported that stacking had halted for about thirty minutes due to some tuning problems.

At 5:59 PM, the NuMI beam permit dropped again. It would not reset. An operator called in an expert.

Operators reported at 6:43 PM that NuMI was back up, but suspected there were still problems.

NuMI experts began a target scan at 8:29 PM.

At 9:32 PM, a NuMI LCW system trip caused the loss of the beam permit. NuMI experts wanted to access the area to investigate. Operators called in the appropriate people to help.

The Recycler began stashing antiprotons at 9:40 PM. There was a problem with the transfers. A Pbar expert investigated and found some bad VSA markers. These markers were corrected, but transfer problems and the investigation continued.

At 11:53 PM, Operations reported that experts were accessing NuMI to investigate the LCW trips. The experts said the LCW system was okay. The problems were due to instrumentation.

Sunday May 8
The midnight shift began with Operations monitoring store 4132, with shots to the Recycler going poorly, and with MiniBooNE taking beam.

The attempts to stash antiprotons ended at 12:32 AM. A Recycler expert continued investigating all the problems. Pbar resumed stacking.

MCenter and NuMI resumed taking beam at 12:43 AM.

At 12:44 AM, BRF5 tripped off. It reset okay.

At 12:50 AM, KRF5 tripped off. It reset okay.

At 12:52 AM, Operations reported three East Anode Power Supply (EAPS) trips. It reset okay each time.

At 12:54 AM, BRF16 tripped off. It reset okay.

At 1:15 AM, EAPS tripped off. Operators investigated and found that they had to turn off and bypass BRF19 before they could get the EAPS back on. This occurred at 2:15 AM.

At 4:54 AM, BRF16 tripped off. It reset okay.

At 6:24 AM, the I- Source chopper tripped off on bad vacuum. Operators had the beam restored ten minutes later.

At 632 AM, BRF16 tripped off. It reset okay.

At 9:37 AM, BRF5 tripped off repeatedly. It did reset.

Operations terminated store 4132 at 12:04 PM.

Around 12:18 PM, EAPS tripped off. It reset okay.

Operators made the Collider Detector at Fermilab (CDF) and D0 ready for access at 12:24 PM.

D0 came out of access at 1:45 PM.

At 1:58 PM, BRF8 tripped off. It reset okay.

At 2:56 PM, BRF11 tripped off. It reset okay.

CDF came out of access around 3 PM.

Operations put the TeV through a dry squeeze at 3:13 PM.

At 3:48 PM, KRF2 tripped off. Operators had it reset five minutes later.

At 4:05 PM, BRF8 tripped off. It reset okay.

Operations put the TeV into shot setup at 4:10 PM.

Operations established store 4134 at 6:35 PM with an initial luminosity of 72.77E30. Pbar resumed stacking.

NuMI and MCenter resumed taking beam at 7:16 PM.

At 7:20 PM, EAPS tripped off. NuMI also tripped off at the same time. Operators reported at 7:38 PM that they had to disable Switchyard (SY) beam, reestablish NuMI beam, and then establish SY beam to get everything back on.

At 8:06 PM, the Tevatron Electron Lens (TEL) tripped off. It reset okay.

At 9:33 PM, Operations reported that according to the NuMI target LCW tank level, the target leak returned. However, the helium backpressure was keeping the system stable.

A Recycler expert began a Boomerang study period at 11:20 PM. (Shots of antiprotons get sent to MI and then sent directly back to the Recycler.)

Monday May 9
The midnight shift began with Operations monitoring store 4134, with Recycler conducting a study, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler expert put the Recycler into a proton only mode at 12:53 AM.

At 1:30 AM, EAPS tripped off. It reset okay.

At 1:48 AM, BRF13 tripped off. It reset okay.

The Recycler began stashing antiprotons at 1:50 AM.

The Recycler stash successfully ended and Pbar resumed stacking at 2:46 AM.

NuMI and MCenter resumed taking beam at 3:01 AM.

At 3:42 AM, Debuncher RF stations 1-7 tripped off. They all reset okay.

At 5:01 AM, the NuMI beam permit tripped off. It reset okay.

At 5:06 AM, MIRF8 tripped off. It reset okay.

At 6:34 AM, KRF7 tripped off. It reset okay.

The Plan for Monday
The plan for today is to terminate store 4134 this afternoon and then establish a combination store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/9/2005   email Fermilab